| <p>On 22/11/2018 in the afternoon:</p>
<p>We increased the ADC thresholds to 40 (previously 8).</p>
<p>For an alpha source with pulser connected, the event multiplicity dropped from ~63 to ~10. Dead time ~ 7%</p>
<p>Checking the timing signals for the near S2:</p>
<ul>
<li>RAL X channels 0 - 5 bad, 6 and 7 good -> disconnected timing</li>
<li>RAL U, XII, XIII, R, and S good</li>
<li>RAL W channels 0 - 5 good, 6 and 7 bad</li>
<li>RAL Q all bad -> disconnected timing</li>
</ul>
<p>Started alpha source run with pulser (same pulser settings as run before)</p>
<ul>
<li>Run # 181122161955</li>
</ul>
<p>After run the Far S2 RAL thresholds set to 10mV, the close S2 thresholds set to 50mV</p>
<p>Started alpha source run, no pulser</p>
<ul>
<li>Run # 181122170947</li>
</ul>
<p>Stopped run due to poor alpha rate -> ECL-Nim ribbon cable disconnected.</p>
<p>Checking bias:</p>
<ul>
<li>Far S2 bias = -130V, leak I = 1.91uA</li>
<li>Close S2 bias = -129.9V, leak I = 13.11uA</li>
<li>diode bias = -70V, leak I = 0uA</li>
</ul>
<p>Vented chamber and checked 12um mylar foil in front of close S2 - no visual damage. LEft</p>
<p>Close detector changed: S2 2623-26 (505um) replaced with S2 2623-18 (508um). Mounted new S2 in close position with 12um mylar</p>
<p>Removed alpha source - needed by other experiment. Otherwise target ladder configuration is same as eLog entry #47</p>
<p>Checked grounding cables going from preamps to ICU boxes were tightly connected</p>
<p>Attempted to bias close S2 -> no leak current! Left lab for the day.</p>
<p>The next morning (23/11/2018):</p>
<p>We found the BNC - SMC bulkhead connector attached to the near S2 preamp box is missing a pin. We dismounted the near S2.</p>
<p>We observed ~50kHz noise on Far S2, this is removed when we lifted the preamp - ICU box ribbon cables away from the chamber vacuum gauges.</p> |